OpenStudy (anonymous):

Whats the answer to: Simplify using i 3+2i/6+5i

sam (.sam.):

Multiply by Conjugate base 6-5i

OpenStudy (anonymous):

To get rid of i on the bottom multiply top and bottom by conjugate

OpenStudy (anonymous):

I just need the answer

mathslover (mathslover):

\[\frac{3+2i}{6+5i}\] \[\frac{(3+2i)(6-5i)}{(6+5i)(6-5i)}\] can u go through out this @ashleyterk remember that we dont answer u directly http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Well, work it out with the information just given to you, it is not so hard to do it....

OpenStudy (anonymous):

@mathslover what do i do next after that?

OpenStudy (anonymous):

FOIL?

OpenStudy (lgbasallote):

remember these concepts: \[(a+bi)(a-bi) = a^2 + b^2\]

OpenStudy (anonymous):

i cant figure it out :(